# London WC1E 6BT, United Kingdom

### v1

@@ Appendix A Calculation of decay and scattering rates
In this appendix, we will give analytical expressions for the scattering cross-sections involving active neutrinos, HNL and ALP in all possible way. Due to the couplings presented in Eq.(9) one can also consider the axion-HNL scattering in early universe. The process happens via t-channel scattering. The matrix element responsible for the process $a(p_{1})+a(p_{2})\rightarrow N(p_{3},s)+N(p_{4},r)$ can be written as
@@
$+\frac{2(s-2m_{a}^{2})\rm{coth}^{-1}\left(\frac{s-2m_{a}^{2}}{\sqrt{s-4m_{a}^{2}}\sqrt{s-4m_{N}^{2}}}\right)}{\sqrt{s-4m_{a}^{2}}\sqrt{s-4m_{N}^{2}}}\Bigg{]}.$ (48)
@@
$+\frac{2(s-2m_{a}^{2})\rm{coth}^{-1}\left(\frac{s-2m_{a}^{2}}{\sqrt{s-4m_{a}^{2}}\sqrt{s-4m_{N}^{2}}}\right)}{\sqrt{s-4m_{a}^{2}}}\Bigg{]},$ (49)
@@
$+\frac{2(2(m_{a}^{2}-m_{N}^{2}+m_{\nu}^{2})-s)\rm{coth}^{-1}\left(\frac{2(m_{a}^{2}-m_{N}^{2}+m_{\nu}^{2})-s}{\sqrt{s-4m_{a}^{2}}\sqrt{s-4m_{\nu}^{2}}}\right)}{\sqrt{s-4m_{a}^{2}}\sqrt{s-4m_{\nu}^{2}}}\Bigg{]}.$ (50)
@@
$+\frac{2(2(m_{a}^{2}-m_{N}^{2})-s)\rm{coth}^{-1}\left(\frac{2(m_{a}^{2}-m_{N}^{2})-s}{\sqrt{s-4m_{a}^{2}}\sqrt{s}}\right)}{\sqrt{s-4m_{a}^{2}}\sqrt{s}}\Bigg{]}.$ (51)
@@
$+\frac{4(2m_{a}^{2}-4m_{N}^{2}-s)\rm{coth}^{-1}\left(\frac{2(m_{a}^{2}-m_{N}^{2})-s}{\sqrt{s-4m_{a}^{2}}\sqrt{s}}\right)}{\sqrt{s-4m_{a}^{2}}\sqrt{s}}\Bigg{]}.$ (52)
@@
$+\frac{2(2(m_{a}^{2}+m_{N}^{2})-s)\rm{coth}^{-1}\left(\frac{2(m_{a}^{2}+m_{N}^{2})-s}{\sqrt{s-4m_{a}^{2}}\sqrt{s-4m_{N}^{2}}}\right)}{\sqrt{s-4m_{a}^{2}}\sqrt{s-4m_{N}^{2}}}\Bigg{]}.$ (53)
@@
$\sigma_{aa\nu\nu}=\frac{4m_{N}^{4}\sqrt{s-4m_{a}^{2}}U_{eN}^{8}}{\pi f_{a}^{4}s^{3/2}}\Bigg{[}-2+\frac{2(s-2m_{a}^{2})\rm{coth}^{-1}\left(\frac{s-2m_{a}^{2}}{\sqrt{s-4m_{a}^{2}}\sqrt{s}}\right)}{\sqrt{s-4m_{a}^{2}}\sqrt{s}}\Bigg{]}.$ (54)
@@
$\sigma_{aa\nu N}=\frac{4m_{N}^{4}\sqrt{s-4m_{a}^{2}}U_{eN}^{6}}{\pi f_{a}^{4}s^{3/2}}\Bigg{[}-2+\frac{m_{N}^{2}}{m_{a}^{2}}+\frac{2(s-2m_{a}^{2})\rm{coth}^{-1}\left(\frac{s-2m_{a}^{2}}{\sqrt{s-4m_{a}^{2}}\sqrt{s}}\right)}{\sqrt{s-4m_{a}^{2}}\sqrt{s}}\Bigg{]}.$ (55)
@@
$+\frac{2(s+2m_{a}^{2}-2m_{N}^{2})\rm{coth}^{-1}\left(\frac{-2m_{a}^{2}+2m_{N}^{2}-s}{\sqrt{s-4m_{N}^{2}}\sqrt{s}}\right)}{\sqrt{s-4m_{N}^{2}}\sqrt{s}}\Bigg{]}.$ (56)
@@
$+m_{N}^{2}(3+4s))\Big{\}}\rm{coth}^{-1}\Bigg{(}\frac{s-2m_{a}^{2}}{\sqrt{s-4m_{a}^{2}}\sqrt{s-4m_{N}^{2}}}\Bigg{)}\Bigg{]}.$ (57)

### v2

@@ Appendix A Calculation of decay and scattering rates
In this appendix, we will give analytical expressions for the scattering cross-sections involving active neutrinos, HNL and ALP in all possible way. Due to the couplings presented in Eq.(2.9) one can also consider the axion-HNL scattering in early universe. The process happens via t-channel scattering. The matrix element responsible for the process $a(p_{1})+a(p_{2})\rightarrow N(p_{3},s)+N(p_{4},r)$ can be written as
@@
$+\frac{2(s-2m_{a}^{2})\rm{coth}^{-1}\left(\frac{s-2m_{a}^{2}}{\sqrt{s-4m_{a}^{2}}\sqrt{s-4m_{N}^{2}}}\right)}{\sqrt{s-4m_{a}^{2}}\sqrt{s-4m_{N}^{2}}}\Bigg{]}.$ (A.3)
@@
$+\frac{2(s-2m_{a}^{2})\rm{coth}^{-1}\left(\frac{s-2m_{a}^{2}}{\sqrt{s-4m_{a}^{2}}\sqrt{s-4m_{N}^{2}}}\right)}{\sqrt{s-4m_{a}^{2}}}\Bigg{]},$ (A.4)
@@
$+\frac{2(2(m_{a}^{2}-m_{N}^{2}+m_{\nu}^{2})-s)\rm{coth}^{-1}\left(\frac{2(m_{a}^{2}-m_{N}^{2}+m_{\nu}^{2})-s}{\sqrt{s-4m_{a}^{2}}\sqrt{s-4m_{\nu}^{2}}}\right)}{\sqrt{s-4m_{a}^{2}}\sqrt{s-4m_{\nu}^{2}}}\Bigg{]}.$ (A.5)
@@
$+\frac{2(2(m_{a}^{2}-m_{N}^{2})-s)\rm{coth}^{-1}\left(\frac{2(m_{a}^{2}-m_{N}^{2})-s}{\sqrt{s-4m_{a}^{2}}\sqrt{s}}\right)}{\sqrt{s-4m_{a}^{2}}\sqrt{s}}\Bigg{]}.$ (A.6)
@@
$+\frac{4(2m_{a}^{2}-4m_{N}^{2}-s)\rm{coth}^{-1}\left(\frac{2(m_{a}^{2}-m_{N}^{2})-s}{\sqrt{s-4m_{a}^{2}}\sqrt{s}}\right)}{\sqrt{s-4m_{a}^{2}}\sqrt{s}}\Bigg{]}.$ (A.7)
@@
$+\frac{2(2(m_{a}^{2}+m_{N}^{2})-s)\rm{coth}^{-1}\left(\frac{2(m_{a}^{2}+m_{N}^{2})-s}{\sqrt{s-4m_{a}^{2}}\sqrt{s-4m_{N}^{2}}}\right)}{\sqrt{s-4m_{a}^{2}}\sqrt{s-4m_{N}^{2}}}\Bigg{]}.$ (A.8)
@@
$\sigma_{aa\nu\nu}=\frac{4m_{N}^{4}\sqrt{s-4m_{a}^{2}}U_{eN}^{8}}{\pi f_{a}^{4}s^{3/2}}\Bigg{[}-2+\frac{2(s-2m_{a}^{2})\rm{coth}^{-1}\left(\frac{s-2m_{a}^{2}}{\sqrt{s-4m_{a}^{2}}\sqrt{s}}\right)}{\sqrt{s-4m_{a}^{2}}\sqrt{s}}\Bigg{]}.$ (A.9)
@@
$\sigma_{aa\nu N}=\frac{4m_{N}^{4}\sqrt{s-4m_{a}^{2}}U_{eN}^{6}}{\pi f_{a}^{4}s^{3/2}}\Bigg{[}-2+\frac{m_{N}^{2}}{m_{a}^{2}}+\frac{2(s-2m_{a}^{2})\rm{coth}^{-1}\left(\frac{s-2m_{a}^{2}}{\sqrt{s-4m_{a}^{2}}\sqrt{s}}\right)}{\sqrt{s-4m_{a}^{2}}\sqrt{s}}\Bigg{]}.$ (A.10)
@@
$+\frac{2(s+2m_{a}^{2}-2m_{N}^{2})\rm{coth}^{-1}\left(\frac{-2m_{a}^{2}+2m_{N}^{2}-s}{\sqrt{s-4m_{N}^{2}}\sqrt{s}}\right)}{\sqrt{s-4m_{N}^{2}}\sqrt{s}}\Bigg{]}.$ (A.11)
@@
$+m_{N}^{2}(3+4s))\Big{\}}\rm{coth}^{-1}\Bigg{(}\frac{s-2m_{a}^{2}}{\sqrt{s-4m_{a}^{2}}\sqrt{s-4m_{N}^{2}}}\Bigg{)}\Bigg{]}.$ (A.12)